Maricopa's Success scholarship fund receives a gift of $ 110000. The money is invested in stocks, bonds, and CDs. CDs pay 4.75 % interest, bonds pay 5.2 % interest, and stocks pay 11.6 % interest. Maricopa Success invests $ 45000 more in bonds than in CDs. If the annual income from the investments is $ 7892.5 , how much was invested in each account?
___ Stocks
____ bonds
___ cds

Answer by Boreal(15235)

CDs have x invested
bonds have x+45000 invested
stocks have 110000-2x-45000 invested, or 66000-2x
.0475x+.052(x+45000)+.116(66000-2x)=7892.50
.0475x+.052x+2340-.232x+7656=7892.50
-.1325x=-2103.50
x=$15875.47 CDs @4.75%=$754.08
x+45000=$60875.47 bonds @ 5.2%=$3165.52
66000-2x=$34249.06 stocks @ 11.6%=$3972.89
they add to $7892.49